OpenStudy (anonymous):

Choose the answer based on the most efficient method as presented in the lesson. If the first step in the solution of the equation -8 - 7x = -5x - 10 is "add 10," then what should the next step be? add 5x add 7x add 8

OpenStudy (anonymous):

add 8?

OpenStudy (anonymous):

@iGreen.

OpenStudy (haleyelizabeth2017):

Incorrect.

OpenStudy (anonymous):

oh ok thnx

OpenStudy (haleyelizabeth2017):

Just write it out without wording... -8 - 7x = -5x - 10 +10 +10 2-7x=-5x What would you generally do next?

OpenStudy (anonymous):

-8 - 7x = -5x - 10 Adding 10 gives us: 2 - 7x = -5x Now we want to get rid of '-7x' so what do we do next?

OpenStudy (anonymous):

thnx